# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ 7       Update the virtual queues,
$\leq n+\frac{\eta n}{2\varepsilon}(BV+1)^{2}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
Summing the above for $t\in\{T_{0},T_{0}+1,\dots,T_{0}+T-1\}$, we have that,
@@
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\right\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right),$
where for the last inequality, we have used Lemma 13. Rearranging the above and using $\lVert\boldsymbol{Q}(T_{0})\rVert^{2}\leq n(BV+1)^{2}$, and $\lVert\boldsymbol{Q}(T+T_{0})\rVert^{2}\geq 0$ we are done. ∎
Now, we are ready to establish the performance bound of the algorithm.
Theorem 2
For any parameters $T,T_{0}\in\mathbb{N}$, $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
In particular
  1. 1.
    For any $\epsilon>0$, choosing $V=\Theta(1/\epsilon)$, $\varepsilon=\Theta(\epsilon)<1/s$, $\eta=\Theta(\epsilon^{3})$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=1}^{T}\boldsymbol{X}(t)\right\}\right)=O(\epsilon).$
  2. 2.
    In the finite time horizon setting with $T,T_{0}\in\mathbb{N}$, using $\eta=\Theta(1/T)$, $\varepsilon=\Theta(1/T^{1/3})<1/s$, and $V=\Theta(T^{1/3})$, we have
    $\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\frac{\log(T)}{T^{1/3}}\right).$
@@
Consider the problem
[…]
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (36)
$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (37)
Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is $\phi^{\text{opt}}$. Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote the solution of the above problem. Substituting $\boldsymbol{p}^{*}(1-\varepsilon n)+\varepsilon\boldsymbol{1},\boldsymbol{\gamma}^{*}$ in Lemma 17 we have that,
$VT\phi^{\text{opt}}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i}(1-\varepsilon n)-\varepsilon q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-q_{i}p^{*}_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}nq_{i}p^{*}_{i}\mathbb{E}\{Q_{i}(t)\}$
@@
$\leq_{(a)}nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\varepsilon n^{2}T(BV+1)+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (38)
where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (34)-(37)), and Lemma 18. Now, we use the Jensen’s inequality to obtain,
$\phi^{\text{opt}}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
Notice that Lemma 11 holds since the queue updates are the same as in the multi-channel case. Combining above with Lemma 11, we have that,
@@
$\leq_{(a)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{\eta n}{2\varepsilon V}(BV+1)^{2}+\frac{\varepsilon n^{2}}{V}(BV+1)+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Lemma 18. Now, using Jensen’s inequality, we are done. ∎
Comparing the error bounds of Theorem 2 and Theorem 1 (with $m=1$), it can be seen that the bound of Theorem 2 is strictly better. In particular, the second term of the error bound has a $\Theta(n)$ dependence on $n$ in Theorem 2, whereas the corresponding value is $\Theta(n^{3})$ in Theorem 1.
V UCB Algorithm
In this section, we present a UCB-based algorithm for the link selection problem in MAC (Algorithm 3), after which we move on to the analysis. Recall that $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ is the user-channel assignment during time slot $t$. For $t\in[s]$, we set $\boldsymbol{Y}(t)$ arbitrarily such that each user-channel pair is explored exactly once during the interval $[s]$. For each $t>s$, we compute a permutation matrix $\boldsymbol{M}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1. The assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ for the $t>s$ is obtained by discarding the last $s-m$ columns or the last $s-n$ rows of $\boldsymbol{M}(t)$ depending on whether $n$ or $m$ is larger. The algorithm uses parameters $\delta_{s},\delta_{s+1},\dots$ that satisfy $\delta_{t}>0$ for all $t\in\{s,s+1,\dots\}$. For each $i\in[n]$, $j\in[m]$, we define,
@@ V UCB Algorithm
1 For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
@@
5       Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{M}(t)$ by solving
@@
$\boldsymbol{M}(t)=\arg\max_{\boldsymbol{M}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
where $\text{UCB}_{i,j}(t-1)$ is defined in (V).
6       Use the assignment $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$, where $Y_{i,j}(t)=M_{i,j}(t)$ for $(i,j)\in[n]\times[m]$ and receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
@@ 7       Update the queues
We first focus on solving the inner problem (40). Notice that the problem to solve to obtain $\boldsymbol{\gamma}(t)$ is the same as for Algorithm 1. The problem to solve to obtain $\boldsymbol{M}(t)$ has a classic max-weight structure (hence, $\boldsymbol{M}(t)$ is a permutation matrix). Hence, we can use the Hungarian algorithm [49] to obtain the solution.
Now, we move on to the analysis of the Algorithm 3. Since the queue updates and the problem to solve to obtain $\boldsymbol{\gamma}(t)$ are the same as in Algorithm 1, we have the same deterministic queue bound. We formally state this in the following lemma.
@@
Proof:
This follows repeating the argument of Lemma 7. ∎
@@
[…]
where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (V). We have the following lemma.
@@
The proof follows directly by applying the Hoeffding inequality with union bound. See [48] for more details. ∎
@@
[…]
@@
Proof is similar to the proof of Lemma 8 and is given in Appendix B. ∎
@@
$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg{|}\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg{|}\Omega_{t}\right\}\leq-V\phi^{\text{opt}}$ (45)
where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{M}(t)$ for intermediate problem (40) in Algorithm 3 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (42); the last inequality follows from (11). Adding the inequalities (21) and (21) and rearranging, we are done. ∎
@@
Dividing both sides by $V(T-s)$, and using Jensen’s inequality, we are done. ∎
@@
Now, we have the following theorem that establishes the performance bound of Algorithm 3.
Theorem 3
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (47)
where $f_{i,j}(t)$ is defined in (V). In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
  1. 1.
    For all $V>0$,
    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
  2. 2.
    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
  3. 3.
    Fix $\epsilon>0$. Using $V=\Theta(1/\epsilon)$, we have
    $\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)=\mathcal{O}(\epsilon),$
    and $\mathcal{O}$ hides the dependence on all parameters but $\epsilon$.
@@
Combining (3) and (3) and using the Jensen’s inequality, we have (3).
@@
which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$.
For part 3, notice that from part 1,
$\phi^{\text{opt}}-\lim\inf_{T\to\infty}\phi\left(\mathbb{E}\left\{\frac{1}{T}\left(\sum_{t=1}^{T}\boldsymbol{X}(t)\right)\right\}\right)\leq\frac{n}{V}.$
Using $V=\Theta(1/\epsilon)$, we are done. ∎
VI Discussion
VI-A Distributed Implementation
Consider a system where the $n$ users take decisions in the absence of a centralized controller. In each time slot, each user receives as feedback whether each channel was successfully accessed or not (this can be, for instance, via sensing each channel). Notice that $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$ are $\mathcal{H}(t)$-measurable, where $\mathcal{H}(t)$ is defined in (43). Hence, during time slot $t$, if each user has access to the full history $\mathcal{H}(t)$, they can independently implement Algorithm 3 to find $\boldsymbol{M}(t),\boldsymbol{\gamma}(t)$. From the feedback, each user can compute $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$ and hence the full history $\mathcal{H}(t+1)$. Hence, starting from $\mathcal{H}(s+1)$, each user can independently implement Algorithm 3. Notice that this is not possible in Algorithm 1, since the channel assignments are not $\mathcal{H}(t)$-measurable.

### v2

@@ VI-A3 Analysis of the Algorithm
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (76)
Notice that from the optimality of $(\boldsymbol{\gamma}(t+1),\boldsymbol{p}(t+1))$ in (5), we have that for any $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$ and $\boldsymbol{\gamma}\in[0,1]^{n}$
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\ \ \ \ +\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}p_{i}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$
$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}-\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$ (77)
where the inequality follows from Lemma 2-2. Using the above in (18), we have for any $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$ and $\boldsymbol{\gamma}\in[0,1]^{n}$
$\Delta(t)$ $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq n+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\}$ (78)
Now, we add the above for $t\in[T_{0}:T_{0}+T-1]$ to get
@@
$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\}$
$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T_{0}))\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(T+T_{0}))\}$
$\leq_{(a)}nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}-VT\phi(\boldsymbol{\gamma})+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)$ (79)
where (a) follows by $D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{s})\leq\log(1/\varepsilon)$ for $\boldsymbol{p}\in\Delta^{n}$ and $\boldsymbol{s}\in\Delta^{n}_{\varepsilon}$ (Proof follows the same steps as Lemma 4). Rearranging and using $\lVert Q(T_{0})\rVert^{2}\leq n(BV+1)^{2}$ (follows from Corollary (1.1)), we are done. ∎
Now, we are ready to establish Theorem 3.
@@
Substituting $(1-\varepsilon n)\boldsymbol{p}^{*}+\varepsilon,\boldsymbol{\gamma}^{*}$ in Lemma 18, where $\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*}$ is the optimal solution of Problem (66)-(69), we have that,
$VT\phi^{*}-V\sum_{t=T_{0}}^{T+T_{0}-1}\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$
$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma^{*}_{i}-p^{*}_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\varepsilon n\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\}$
@@
$\leq_{(a)}nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\varepsilon n^{2}(BV+1)T+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (80)
where (a) follows since $q_{i}p_{i}^{*}\geq\gamma_{i}^{*}$ (since $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ is feasible for problem (66)-(69)), and Corollary (1.1). Dividing both sides by $VT$ and using Jensen’s inequality, we have
$\phi^{*}-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{\gamma}(t+1)\right)\right\}$
$\leq\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}$ (81)
Combining above with Lemma 11, we have that,
@@
$\leq_{(a)}\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{\sqrt{n}B\mathbb{E}\{\lVert\boldsymbol{Q}(T+T_{0})\rVert\}}{T}$
$\leq_{(b)}\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T},$
where (a) follows due to Lemma 11, and (b) follows due to Corollary (1.1). Now, using Jensen’s inequality, we are done. ∎
VI-B Non-Adaptive MAC
In this section, we present a UCB-based algorithm for the link selection problem in MAC (Algorithm 4), after which we move on to the analysis. For $t\in[s]$, we set $\boldsymbol{Y}(t)$ arbitrarily such that each user-channel pair is explored exactly once during the interval $[s]$. For each $t>s$, we compute permutation matrix $\boldsymbol{Y}(t)$, an auxiliary vector $\boldsymbol{\gamma}(t)$, and a virtual queue $\boldsymbol{Q}(t)=[Q_{1}(t),\dots,Q_{n}(t)]$ similar to Algorithm 1. The algorithm uses parameters $\delta_{s},\delta_{s+1},\dots$ that satisfy $\delta_{t}>0$ for all $t\in\{s,s+1,\dots\}$. For each $i\in[n]$, $j\in[m]$, we define,
@@ VI-B Non-Adaptive MAC
1For $t\in[s]$, choose $\boldsymbol{Y}(t)\in\{0,1\}^{n\times m}$ arbitrarily such that each user-channel link is explored exactly once.
@@
5   Find $\boldsymbol{\gamma}(t)$ and $\boldsymbol{Y}(t)$ by solving
@@
$\boldsymbol{Y}(t)=\arg\max_{\boldsymbol{Y}\in\mathcal{S}^{\text{doub}}}\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\text{UCB}_{i,j}(t-1)Y_{i,j},$
where $\text{UCB}_{i,j}(t-1)$ is defined in (VI-B).
6   Receive feedback $\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
@@ 7   Update the queues
We first state the performance bound of Algorithm 4.
Theorem 4
$\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T},$ (85)
where $f_{i,j}(t)$ is defined in (VI-B) and $\phi^{\text{max}}=\max_{\boldsymbol{\gamma}\in[0,1]^{n}}\phi(\boldsymbol{\gamma})$. In particular using $\delta_{t}=1/t$ for all $t\geq s$, we have
  1. 1.
    For all $V>0$,
    $\ \ \ \ +\frac{nB(BV+1)}{T-s}+\frac{(\sqrt{n}B+1)s}{T}.$
  2. 2.
    Assume $T\geq 2s$. Using $V=\Theta(\sqrt{T})$, we have
    and $\mathcal{O}$ hides the dependence on all parameters but $T$.
VI-B1 Solving Problem (83)
We first focus on solving the inner problem (83). Notice that the problem to solve to obtain $\boldsymbol{\gamma}(t)$ is the same as for Algorithm 1. The problem to solve to obtain $\boldsymbol{Y}(t)$ has a classic max-weight structure (hence, $\boldsymbol{Y}(t)$ is a permutation matrix). Hence, we can use the Hungarian algorithm [32] to obtain the solution.
VI-B2 Analysis of the Algorithm
Now, we focus on proving Theorem 4. Since the queue updates and the problem to solve to obtain $\boldsymbol{\gamma}(t)$ are the same as in Algorithm 1, we have the same deterministic queue bound. We formally state this in the following lemma.
@@ VI-B2 Analysis of the Algorithm
$\Omega_{t}=\Bigg\{$ $\text{UCB}_{i,j}(t-1)-2f_{i,j}(t-1)\leq q_{i,j}\leq\text{UCB}_{i,j}(t-1),\forall i\in[n],j\in[m]\Bigg\},$ (86)
where $\text{UCB}_{i,j}(t-1)$ and $f_{i,j}(t-1)$ are defined in (VI-B). We have the following lemma.
@@
The proof follows directly by applying the Hoeffding inequality with union bound. See [lattimore_szepesvári_2020] for more details. ∎
@@
$\mathcal{H}(t)=\{$ $\boldsymbol{Y}(1),\dots,\boldsymbol{Y}(t-1),\boldsymbol{Y}(1)\odot\boldsymbol{S}(1),\dots,\boldsymbol{Y}(t-1)\odot\boldsymbol{S}(t-1)\},$ (87)
@@
Notice that from the queuing equation (84), we have
$\leq Q^{2}_{i}(t)+2+2Q_{i}(t)[\gamma_{i}(t)-X_{i}(t)]$
for all $i\in[n]$. Summing the above for $i\in[n]$, we have
where the last equality follows since $\boldsymbol{Y}(t),\boldsymbol{\gamma}(t),\boldsymbol{Q}(t)$ are $\mathcal{H}(t)$-measurable. Taking expectations, we have the result. ∎
@@
$\leq_{(b)}\mathbb{E}\left\{-V\phi(\boldsymbol{\gamma}^{*})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}^{*}\Bigg|\Omega_{t}\right\}-\mathbb{E}\left\{\sum_{i=1}^{n}Q_{i}(t)\sum_{j=1}^{m}q_{i,j}P_{i,j}^{*}\Bigg|\Omega_{t}\right\}$
$\leq_{(c)}-V\phi^{*}\leq-V\phi^{\text{opt}}$ (89)
where (a) follows from the optimality of $\boldsymbol{\gamma}(t),\boldsymbol{Y}(t)$ for the intermediate problem (83) in Algorithm 4 (recall $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is the optimal solution of P2); (b) follows from the definition of the good event $\Omega_{t}$ in (86); (c) follows from the constraint (15); the last inequality follows from Lemma 1. Adding the inequalities (22) and (22) and rearranging, we are done. ∎
@@
Dividing both sides by V(Ts)V(T-s), and using Jensen’s inequality, we are done. ∎
@@
Now, we are ready to prove Theorem 4.
@@
Combining (VI-B2) and (VI-B2) and using the Jensen’s inequality, we have (4).
@@
which is clearly $\mathcal{O}\left(\sqrt{\frac{\log(T)}{T}}\right)$ when $V=\Theta(\sqrt{T})$. ∎
VI-C Single-Channel Case with Minimum Utility
Consider the single-channel case $m=1$ with utility function $\phi(\boldsymbol{x})=\min\{x_{1},\dots,x_{n}\}$. Observe that if $q_{i}=0$ for some $i\in[n]$, then $\phi\left(\mathbb{E}\{\boldsymbol{\bar{X}}(T)\}\right)=0$ under any assignment policy. Consequently, $\phi^{\text{opt}}=0$. Therefore, without loss of generality, we assume that $q_{i}>0$ for all $i\in[n]$.
If the values $q_{1},\dots,q_{n}$ are known, an optimal policy for (P1) is to sample user $i$ with probability $p_{i}^{*}=\frac{1/q_{i}}{\sum_{k=1}^{n}(1/q_{k})}$. Under this policy, $\mathbb{E}[X_{i}(t)]=q_{i}p_{i}^{*}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}$ for all $i\in[n]$. Hence, we have $\phi^{\mathrm{opt}}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}$.
Now consider the following procedure:
  • Sample a user uniformly at random.
  • Continue to serve the user until a successful transmission.
  • Resample a user uniformly at random and repeat.
This mechanism does not require knowledge of the $q_{i}$. Fix a user $i\in[n]$. Let $(T_{j},R_{j})_{j\geq 1}$ denote the renewal process where, $T_{j}$ is the number of slots required until the $j$-th sampled user achieves a success, $R_{j}=1$ if the $j$-th sampled user is $i$, and $R_{j}=0$ otherwise. Note that the sequence ${(T_{j},R_{j})}_{j\geq 1}$ is independent and identically distributed. Since users are sampled uniformly at random, and conditional on sampling user $k$, the transmission duration is geometrically distributed with mean $1/q_{k}$, we have $\mathbb{E}\{R_{1}\}=\frac{1}{n}$ and $\mathbb{E}[T_{1}]=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{q_{k}}$. By the renewal reward theorem, with probability 1,
$\lim_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}X_{i}(t)=\frac{\mathbb{E}[R_{1}]}{\mathbb{E}[T_{1}]}=\frac{1}{\sum_{k=1}^{n}(1/q_{k})}=\phi^{\text{opt}}.$ (96)
Thus every user achieves throughput $\phi^{\text{opt}}$, and therefore
$\lim_{T\to\infty}\min\{\bar{X}_{1}(T),\dots,\bar{X}_{n}(T)\}=\phi^{\mathrm{opt}}$ (97)
with probability 1. Hence, the mechanism converges to the optimal value without requiring knowledge of $q_{1},\dots,q_{n}$.
Suppose now that the values $q_{i}$ change at some time. Because geometric service times are memoryless, the residual service time at the change point has the same distribution as a fresh geometric random variable with the new parameter. Therefore, from the change point onward, the system evolves as a delayed renewal process under the new parameters, and the renewal argument still applies. Hence, the process is adaptive.